OpenStudy (anonymous):

Choose the function whose graph is given by (graph will be inside!) InPool=Pool1 A. y = 3sin(x - 2) + 1 B. y = 6sin(x - 2) - 2 C. y = 3cos(x - 3) + 1 D. y = 3sin(x - 2) + 2 ***im not sure :( please explain?

OpenStudy (anonymous):

graph!

OpenStudy (mertsj):

It goes from a low of -2 to a high of 4 so the centerline is (-2+4)/2=1

OpenStudy (mertsj):

So it has to be either a or c

OpenStudy (mathstudent55):

A plain sine curve y = sin(x) goes from y = -1 to y = 1. That means an amplitude of 2. When you modify it by multiplying by a number such as y = a*sin x, you increase its amplitude to 2a.

OpenStudy (anonymous):

oh okay :) so if it's amplitude of 2, would the answer be this then? A. y = 3sin(x - 2) + 1 ?? is that it? ;/

OpenStudy (mathstudent55):

Since this graph goes from y = -2 to y = 4, the amplitude is 6, so a = 3. You need y = 3*sin(x) But now you also have to deal with the fact that it is also shifted horizontally and vertically.

OpenStudy (mertsj):

It's amplitude is 3.

OpenStudy (anonymous):

okay, so because of that, the answer is A. y = 3sin(x - 2) + 1 right? and amplitude is the part b4 sin right?

OpenStudy (mathstudent55):

Since instead of going from y = -3 to y = 3, it goes from y = -2 to y = 4, that means it has a vertical shift of +1. Thar means y = sin(x) + k, where k = 1 So now you already have y = 3*sin(x) + 1 Now you need to account for the horizontal shift.

OpenStudy (mertsj):

So it all comes down to whether it is a sin that has been moved right 2 or a cosine that has been moved right 3

OpenStudy (anonymous):

okay :) how do i do that? :/

OpenStudy (anonymous):

find the horizontal shift i mean?

OpenStudy (anonymous):

is the horizontal shift 2?? because of this? y = -2 to y = 4

OpenStudy (mertsj):

What you need to do is draw the centerline which is y = 1. Then you will be able to determine the horizontal displacement.

OpenStudy (anonymous):

okay, there's an equation i need to follow right?

OpenStudy (mertsj):

It appears to me that it is a sine that has been moved 2 units to the right.

OpenStudy (mertsj):

No. There is not a formula for that. I think you are talking about the formula for finding the period.

OpenStudy (anonymous):

okay, so answer A? y = 3sin(x - 2) + 1 because amplitude matches: 3 and moves to the right 2 units: -2 :) did i get that? and ohhh yeah! the period! my bad!!

OpenStudy (mertsj):

yes

OpenStudy (anonymous):

awesome! :) yayy :D thanks!! @Mertsj can u pls give @mathstudent55 a medal? i can only give one :(

OpenStudy (mathstudent55):

Yes, I agree with @Mertsj When you bring down the curve 1 unit, the point that is currently (2, 1) becomes (2, 0). Since sin(x) goes through (0, 0), and this curve goes through (2, 0), the horizontal shift is 2 to the right. That is accounted for in y = sin(x - h), where h is the horzontal shift. So putting it all together you get y = 3*sin(x - 2) + 1

OpenStudy (anonymous):

thanks guys :)

OpenStudy (mathstudent55):

wlcm

OpenStudy (mathstudent55):

wlcm